Use the following information to answer the questions below. Assume that the capital account is equal to zero.

Exports of goods and services

540

Primary income received

20

Secondary income received

380

Imports of goods and services

270

Primary income paid abroad

180

Secondary income paid

10

Net acquisition of financial assets

700

Net incurrence of liabilities

80

Net change in financial derivatives

100

What is the trade​ balance?

​(Enter your response as a whole​ number.)What is the current account​ balance? 

​ (Enter your response as a whole​ number.)What is the financial account​ balance?  

​(Enter your response as a whole​ number.)What is the statistical​ discrepancy?  

​(Enter your response as a whole​ number.)
Answer #1

Trade Balance= Export-Import= 540-270 =270

Current Account Balance = Trade balance+ Net primary income+ Net secondary income

Current Account Balance= (540-270)+(20-180)+(380-10)= 270-160+370= 480

Financial Account Balance= Net acquisition of financial assets- net incurrence of liabilities+ Net change in financial derivatives

Financial account balance= 700-80+100= 720

Statistical discrepancy= Financial account balance-(Current Account Balance + Capital account balance)

Statistical discrepancy= 720-(480+0)= 240